OpenStudy (anonymous):

Could someone help me? f(x)=2x+1 and g(x)=-5x+2 when x=3

OpenStudy (anonymous):

f(3)=7

OpenStudy (supernova_sonntag):

f(3)=2(3)+1 Write out your equation with 3 as x. f(3)=6+1 Solve. f(3)=7 ===================================================== g(3)=-5(3)+2 Write out your equation with 3 as x. g(3)=-15+2 Solve. g(3)= -13